B. Crepieux, G. Focker, J. Lettry, M. Lindroos, S. Meunier, A. Muller, T. Nilsson, J. Parra-Lopez, F. Wenander
ISOLDE at URL: http://isolde.cern.ch
Minutes at URL: http://ps-div.web.cern.ch/ps-div/Meetings/ISOLDE/TechMtgs.html
Consolidation projects at URL: http://isolde-consolidation.web.cern.ch/isolde-consolidation/
PSMB meetings at URL: http://cern.web.cern.ch/CERN/Divisions/PS/Meetings/PSMB.html
- The UC target was operating successfully last week for collections.
- A vacuum fault in the front-end region stopped the run prematurely on Saturday. It proved to be one of the two turbo pumps on the front-end itself that had broken. Normally, the front end can operate with only one pump but as no vacuum specialist could be found in the weekend, to do the necessary manipulations to enable this, the run had to be stopped.
- The broken pump is still under guarantee and will be changed in due course. The plan is to operate the GPS with only one turbo pump. The front-ends have been designed for this with one of the turbo pumps acting as reserve for the other. The only consequence will be that the heating (and pumping) of the target will take longer time. On Friday this week the operation will be tried with the next GPS target (a Sn target). In the case that it proves unsatisfactory to complete the 2002 run in this fasion a decision will have to be made by the divisional management on when an intervention to change the pump can be authorized. A spare pump exists and the change itself will require an hours access to the front-end.
- The MgO target was mounted on the separator on Wednesday, a bottle containing Xe-Ne-Ar-Kr gas was mounted for the gas leak as the experiment request Ne.
- During the setting-up the faraday cup HRS.FC300 blocked in the in-position. Gerrit Jan has taken it out and has disconnected the control. At the same intervention it was found that HRS.FC690 had been disconnected. It was reconnected and tested (OK).
- It is now possible to register, submit abstracts or only pre-register on the SAFERIB web site (http://cern.ch/saferib).
- Note that the web site for the technical minutes have changed: http://ps-div.web.cern.ch/ps-div/Meetings/ISOLDE/TechMtgs.html
- The shipping service will in the future only be able to ship radioactive sources in the morning. To avoid unnecessary delays, please, make requests for shipping in advance.
D. Carminati from 17.07 to 30.07
G. Focker from 20.07 to 18.08
M. Lindroos from 16.07 to 05.08 (replaced by J. Lettry from 16.07 to 21.07 and by R. Catherall from 22.07 to 05.08)
J. Lettry from 22.07 to 10.08
T. Nilsson from 11.07 to 31.07 (replaced by F. Herfurth from 12.07 to 14.07 and from 20.07 to 28.07, replaced by Marc Dietrich from 15.07 to 19.07 and from 29.07 to 31.07)
Tuesday 16 July 2002 at 11.00 in the EP Conference Room A, building 304-1-001.